# In-Context Learning with Iterative Demonstration Selection

### v3

@@ 1 Introduction
Despite the effectiveness, the performance of ICL has been shown to be highly sensitive to the selection of demonstration examples (Zhao et al., 2021). Different sets of demonstrations can yield performance ranging from nearly random to comparable with state-of-the-art models (Gao et al., 2021; Lu et al., 2022). To alleviate the above issue, researchers in ICL have proposed a number of methods to select a set of examples as few-shot demonstrations (Rubin et al., 2022; Liu et al., 2022; Li and Qiu, 2023; Wang et al., 2023b; Li et al., 2023a; Ma et al., 2023; An et al., 2023b). However, for LLMs for which parameters or detailed output distributions are not available (Sun et al., 2022), it is still a common practice to randomly select examples or select examples that are semantically similar to the test sample as demonstrations, i.e., considering diversity or similarity. While several approaches investigate the combination of similarity and diversity when prompting with explanations, exploring compositional generalization, or choosing examples for annotation (Ye et al., 2023b; An et al., 2023a; Su et al., 2023), it is not yet clear how to determine and leverage the optimal dimension for different tasks in ICL and how the rationale for answering the query benefits the balance between these two dimensions.
@@ 2.1 Few-shot Learning
Few-shot learning aims to learn tasks with only a few labeled samples, which results in a big challenge, i.e., over-fitting, for models as they typically require large amounts of data for training. Prior methods to address over-fitting mainly focused on augmenting the few-shot data (Gao et al., 2020; Qin and Joty, 2022), reducing the hypothesis space (Triantafillou et al., 2017; Hu et al., 2018), or optimizing the strategy for searching the best hypothesis (Ravi and Larochelle, 2017; Finn et al., 2017). More recently, LLMs have demonstrated strong few-shot learning ability through in-context learning without any parameter updates (Brown et al., 2020).
@@ 2.4 Chain-of-Thought Reasoning
Chain-of-thought (CoT) reasoning induces LLMs to produce intermediate reasoning steps before generating the final answer (Wei et al., 2022b). Depending on whether there are manually designed demonstrations, current CoT reasoning methods mainly include Manual-CoT and Zero-shot-CoT. In Manual-CoT, human-labeled reasoning paths are used to perform CoT reasoning (Wei et al., 2022b; Zhou et al., 2022; Wang et al., 2022b; Li et al., 2022; Wang et al., 2022a). In contrast, LLMs leverage self-generated rationales for reasoning in Zero-shot-CoT (Kojima et al., 2022; Zelikman et al., 2022; Zhang et al., 2023a; Diao et al., 2023). The ongoing research on CoT reasoning has also explored (i) multimodal reasoning (Zhang et al., 2023b; Wu et al., 2023), (ii) distilling knowledge from LLMs (Ho et al., 2022; Fu et al., 2023), and (iii) iterative optimization (Shinn et al., 2023; Madaan et al., 2023; Paul et al., 2023).
@@ 6.3 Analysis
In addition, we show the results with different numbers of demonstrations, the robustness of IDS to different embedding models and Zero-shot-CoT triggers, and the results on two additional datasets in Appendix A.7 $\sim$ A.10, respectively.
@@ A.9 Robustness to Zero-shot-CoT Triggers
A.10 Two Additional Datasets
MNLI Emotion
Top-$k$-Consistency 65.7 58.1
IDS 67.4 60.3
Table 14: Results on two additional datasets.
To better demonstrate the generalization ability of IDS, we further conduct experiments on two additional datasets: MNLI (natural language inference) (Williams et al., 2018) and Emotion (emotion classification) (Saravia et al., 2018). The comparison between IDS and the baseline Top-$k$-Consistency is shown in Table 14, which verifies the strong generalizability of IDS.

### v4

@@ 1 Introduction
Despite the effectiveness, the performance of ICL has been shown to be highly sensitive to the selection of demonstration examples (Zhao et al., 2021). Different sets of demonstrations can yield performance ranging from nearly random to comparable with state-of-the-art models (Gao et al., 2021; Lu et al., 2022; Qin et al., 2023b). To alleviate the above issue, researchers in ICL have proposed a number of methods to select a set of examples as few-shot demonstrations (Rubin et al., 2022; Liu et al., 2022; Li and Qiu, 2023; Wang et al., 2023b; Li et al., 2023a; Ma et al., 2023; An et al., 2023b). However, for LLMs for which parameters or detailed output distributions are not available (Sun et al., 2022), it is still a common practice to randomly select examples or select examples that are semantically similar to the test sample as demonstrations, i.e., considering diversity or similarity. While several approaches investigate the combination of similarity and diversity when prompting with explanations, exploring compositional generalization, or choosing examples for annotation (Ye et al., 2023b; An et al., 2023a; Su et al., 2023), it is not yet clear how to determine and leverage the optimal dimension for different tasks in ICL and how the rationale for answering the query benefits the balance between these two dimensions.
@@ 2.1 Few-shot Learning
Few-shot learning aims to learn tasks with only a few labeled samples, which results in a big challenge, i.e., over-fitting, for models as they typically require large amounts of data for training. Prior methods to address over-fitting mainly focused on augmenting the few-shot data (Gao et al., 2020; Qin and Joty, 2022a, b), reducing the hypothesis space (Triantafillou et al., 2017; Hu et al., 2018; Qin et al., 2023a), or optimizing the strategy for searching the best hypothesis (Ravi and Larochelle, 2017; Finn et al., 2017; Xia et al., 2024). More recently, LLMs have demonstrated strong few-shot learning ability through in-context learning without any parameter updates (Brown et al., 2020).
@@ 2.4 Chain-of-Thought Reasoning
Chain-of-thought (CoT) reasoning induces LLMs to produce intermediate reasoning steps before generating the final answer (Wei et al., 2022b). Depending on whether there are manually designed demonstrations, current CoT reasoning methods mainly include Manual-CoT and Zero-shot-CoT. In Manual-CoT, human-labeled reasoning paths are used to perform CoT reasoning (Wei et al., 2022b; Zhou et al., 2022; Wang et al., 2022b; Li et al., 2022; Wang et al., 2022a; Zhao et al., 2023). In contrast, LLMs leverage self-generated rationales for reasoning in Zero-shot-CoT (Kojima et al., 2022; Zelikman et al., 2022; Zhang et al., 2023a; Diao et al., 2023; Qin et al., 2024). The ongoing research on CoT reasoning has also explored (i) multimodal reasoning (Zhang et al., 2023b; Wu et al., 2023), (ii) distilling knowledge from LLMs (Ho et al., 2022; Fu et al., 2023), and (iii) iterative optimization (Shinn et al., 2023; Madaan et al., 2023; Paul et al., 2023; Jiao et al., 2024).
@@ 6.3 Analysis
In addition, we show the results with different numbers of demonstrations, the robustness of IDS to different embedding models and Zero-shot-CoT triggers in Appendix A.7 $\sim$ A.9, respectively.